# Interpretable Hybrid-Rule Temporal Point Processes

### v2

@@ V-A Experimental Setup
V-B Performance Analysis
In this section, we compare the performance of model fitting, event occurrence time prediction, and rule accuracy using three evaluation metrics. The results shown in Table II demonstrate that HRTPP outperforms baseline models. The lowest NLL indicates that HRTPP effectively captures and models uncertainty. HRTPP also achieves the lowest RMSE, suggesting precise event time predictions and enhancing reliability in medical applications. High rule accuracy of HRTPP demonstrates the extraction of medically relevant rules, improving interpretability and supporting transparent decision-making.
TABLE II: Comparisons of main results.
@@ V-B Performance Analysis
V-C Rule Analysis
In this section, we analyze the quality of the rules from the perspectives of rule expression, correctness, reliability, and distribution, thereby verifying the interpretability of the model. We take the CAD dataset as an example. Fig. III shows top 10 rules of four models. The full rule results for the four datasets are shown in the Appendix.
V-C1 Rule Expression Analysis
In the CAD dataset, we set mortality as the target variable and compare HRTPP with baseline models. We refer to medical guidelines [29] to determine the correctness of the rules. Table III shows the rules, weights, and correstness. From the experimental results, TELLER and CLNN rely on simple variables and discover few rules, missing complex physiological interactions. CLUSTER offers more flexible rules but is highly sensitive to noise, making it difficult to extract correct medical indicators. In contrast, HRTPP uncovers clinically meaningful rules, such as ’if a high anion gap occurs before low heart rate, then death occurs’, highlighting the link between metabolic disorders and circulatory failure. These rules enhance accuracy and align with medical logic, aiding clinical decision-making.
@@ V-C1 Rule Expression Analysis
TABLE III: Rules on the CAD dataset.
Note: The predicates in green boxes are direct indicators, and those in blue boxes are indirect indicators. Model Rule Weight Correctness TELLER Respiratory Rate High → Dead 0.2859 Correct Arterial Blood Pressure diastolic Low → Dead 0.2202 Correct Heart Rate High → Dead 0.1614 Correct Arterial Blood Pressure mean Low → Dead 0.1400 Correct O2 saturation pulseoxymetry Low → Dead 0.1283 Incorrect CLNN c_Glucose High - c_Dead > -0.25 1.03 Correct c_BUN High - c_Dead > -0.63 1.01 Correct c_INR Low - c_BUN High > 0.01 0.99 Incorrect c_Glucose High - c_AST Low > -0.08 0.99 Incorrect c_Albumin High - c_Glucose High > 0.02 0.97 Correct c_BUN High - c_C-Reactive-Protein High > -0.14 0.96 Correct c_Brain Natiuretic Peptide High - c_BUN High > 0.41 0.94 Correct CLUSTER BUN Low equal Lactic Acid Low - Incorrect Hemoglobin High equal Hematocrit High - Incorrect Brain Natiuretic Peptide High after ALT Low - Correct Heart Rate High before AST Low - Correct Calcium ionized High equal Total Bilirubin High - Incorrect CK-MB fraction High equal Calcium ionized Low - Correct Calcium ionized Low equal Cholesterol High - Correct Temperature Low equal Hemoglobin High - Correct Hematocrit Low equal Total Bilirubin Low - Incorrect Glucose Low before Anion gap Low - Incorrect HRTPP Heart Rate Low equal Arterial Blood Pressure diastolic High → Dead 2.4073 Correct Anion gap High before Heart Rate Low → Dead 1.7742 Correct Heart Rate Low equal Potassium High → Dead 1.6257 Correct Potassium Low before O2 saturation pulseoxymetry Low → Dead 1.3959 Correct Respiratory Rate Low before Potassium High → Dead 1.1533 Correct Arterial Blood Pressure mean Low before Arterial Blood Pressure diastolic Low → Dead 1.1312 Correct Anion gap Low equal O2 saturation pulseoxymetry Low → Dead 1.1303 Correct Glucose Low equal Arterial Blood Pressure systolic Low → Dead 0.5603 Correct Heart Rate Low equal INR High → Dead 0.4213 Correct Temperature High and Respiratory Rate High → Dead 0.3194 Correct
V-C2 Rule Distribution Analysis
Understanding mortality risk indicators in CAD patients is crucial for better management and prediction. To analyze these factors, we created a graph network based on medical guidelines [29], visually mapping their hierarchical relationships in Fig. 2. As shown in the Fig. 2, arterial blood pressure, heart rate, C-reactive protein and cholesterol directly effect mortality, while other indicators influence mortality through them. Hence, we classify the rules into direct or indirect rules according to their relationships. Table III shows direct indicators in green and indirect indicators in blue. HRTPP identifies the most direct rules, showing strong ability in capturing key CAD mortality factors. HRTPP also balances direct and indirect rules well, focusing on major factors without omitting minor details, thus enhancing its applicability in medical decision-making.
V-D Stability Analysis
Model stability is crucial in medical decision support, ensuring consistency across conditions. To evaluate model stability, we recorded the frequency of four direct CAD indicators in five runs. Fig. 3 presents the results, where HR, ABP, Cho, and CRP denote Heart Rate, Arterial Blood Pressure, Cholesterol, and C-reactive Protein, respectively. As shown in the figure, HRTPP has the highest frequency of direct rules, which shows the best stability and consistency. Clinically, its stability ensures interpretable results by minimizing fluctuations caused by minor data changes, thereby enhancing decision-making accuracy and supporting personalized treatment.
HRABPChoCRPAvg.00%$20$%$40$%$60$%$80$%$100$%$80$%$100$%$40$%$20$%$80$%$40$%$60$%$20$%00%$30$%$60$%$100$%$60$%$80$%$75$%$100$%$100$%$80$%$60$%$85$%Medical indicationsFrequencyTELLERCLNNCLUSTERHRTPP
@@ V-D Stability Analysis
V-E Case Study
In this section, we show the mechanism by which rules and events affect the target. When a rule is triggered, its strength increases instantaneously, reflecting the potential risk associated with the condition. Fig. 4 shows the intensity curve of a CAD patient. In this figure, abnormalities in physiological indicators triggered key rules over time. At 17.19h, high blood glucose and creatinine appeared, followed by infection-related signs, increasing rule strength. At 60.26h, metabolic imbalance was triggered by an elevated anion gap and low heart rate. At 92.43h, high INR and low heart rate indicated coagulation abnormalities. Finally, at 123.58h, decreased oxygen saturation and increased blood pressure triggered the end-event rule, leading to a sharp rise in intensity and death.
Figure 4: Intensity over time for a CAD patient. The red curve represents the intensity of the target variable, reflecting the probability of death. Key events affecting the mortality are annotated below the curve, while the timing of logic rules is marked above. The entire figure illustrates the patient’s full clinical progression in the ICU.
This mechanism enables real-time monitoring and early warning of high-risk conditions while providing interpretable analysis of disease progression to support clinical decisions. By applying rule-guided dynamic analysis, we identify key causal relationships in CAD progression and detect turning points in disease deterioration, allowing clinicians to intervene precisely at critical moments. These rules are not just data-driven predictions but verifiable medical reasoning, ensuring transparency and reliability in decision-making while improving the efficiency and accuracy of CAD diagnosis and treatment.
V-F Ablation Study
We conducted an ablation study on four datasets to assess the impact of Numerical Feature Augmentation (NFA) on model performance. We classify models into two categories. Category I uses only basic intensity, like probabilistic TPP models. Category II uses basic and rule-based intensity. Table IV shows the NLL of two categories. The experimental results indicate that incorporating NFA significantly improves model performance. These findings validate the effectiveness of numerical intensity in medical event modeling.
TABLE IV: Ablation study of numerical feature augmentation.

### v3

@@ 5.1 Experimental Setup
5.2 Performance Analysis
In this section, we compare the performance of model fitting, event occurrence time prediction, and rule accuracy using three evaluation metrics. The results shown in Table 2 demonstrate that HRTPP outperforms baseline models. The lowest NLL indicates that HRTPP effectively captures and models uncertainty. HRTPP also achieves the lowest RMSE, suggesting precise event time predictions and enhancing reliability in medical applications. High rule accuracy of HRTPP demonstrates the extraction of medically relevant rules, improving interpretability and supporting transparent decision-making. Among the two most accurate models, HRTPP is computationally more efficient, consistently completing training in under 16 hours on all datasets, whereas TELLER requires over 48 hours.
Table 2: Comparisons of main results.
@@ 5.2 Performance Analysis
5.3 Rule Analysis
In this section, we analyze the quality of the rules from the perspectives of rule expression, correctness, reliability, and distribution, thereby verifying the interpretability of the model. We take the CAD dataset as an example. Table 3 shows top 10 rules of four models. The full results are shown in the supplementary material.
Table 3: Rules on the CAD dataset. The predicates in green boxes are direct indicators, and those in blue boxes are indirect indicators.
Model Rule Weight Correctness
TELLER Respiratory Rate High → Dead 0.2859 Correct
Arterial Blood Pressure diastolic Low → Dead 0.2202 Correct
Heart Rate High → Dead 0.1614 Correct
Arterial Blood Pressure mean Low → Dead 0.1400 Correct
O2 saturation pulseoxymetry Low → Dead 0.1283 Incorrect
CLNN c_Glucose High - c_Dead > -0.25 1.03 Correct
c_BUN High - c_Dead > -0.63 1.01 Correct
c_INR Low - c_BUN High > 0.01 0.99 Incorrect
c_Glucose High - c_AST Low > -0.08 0.99 Incorrect
c_Albumin High - c_Glucose High > 0.02 0.97 Correct
c_BUN High - c_C-Reactive-Protein High > -0.14 0.96 Correct
c_Brain Natiuretic Peptide High - c_BUN High > 0.41 0.94 Correct
CLUSTER BUN Low equal Lactic Acid Low - Incorrect
Hemoglobin High equal Hematocrit High - Incorrect
Brain Natiuretic Peptide High after ALT Low - Correct
Heart Rate High before AST Low - Correct
Calcium ionized High equal Total Bilirubin High - Incorrect
CK-MB fraction High equal Calcium ionized Low - Correct
Calcium ionized Low equal Cholesterol High - Correct
Temperature Low equal Hemoglobin High - Correct
Hematocrit Low equal Total Bilirubin Low - Incorrect
Glucose Low before Anion gap Low - Incorrect
HRTPP Heart Rate Low equal Arterial Blood Pressure diastolic High → Dead 2.4073 Correct
Anion gap High before Heart Rate Low → Dead 1.7742 Correct
Heart Rate Low equal Potassium High → Dead 1.6257 Correct
Potassium Low before O2 saturation pulseoxymetry Low → Dead 1.3959 Correct
Respiratory Rate Low before Potassium High → Dead 1.1533 Correct
Arterial Blood Pressure mean Low before Arterial Blood Pressure diastolic Low → Dead 1.1312 Correct
Anion gap Low equal O2 saturation pulseoxymetry Low → Dead 1.1303 Correct
Glucose Low equal Arterial Blood Pressure systolic Low → Dead 0.5603 Correct
Heart Rate Low equal INR High → Dead 0.4213 Correct
Temperature High and Respiratory Rate High → Dead 0.3194 Correct
5.3.1 Rule Expression Analysis.
In the CAD dataset, we set mortality as the target variable and compare HRTPP with baseline models. The correctness of the rules was determined by referencing medical guidelines jointly issued by authoritative medical bodies [22]. Table 3 shows the rules, weights, and correctness. The weights of HRTPP are $\alpha_{j}$ in Eq.(8), demonstrating the importance of rules. Different rules are connected by logical disjunction relationships. From the experimental results, TELLER and CLNN rely on simple variables and discover few rules, missing complex physiological interactions. CLUSTER offers more flexible rules but is highly sensitive to noise, making it difficult to extract correct medical indicators. In contrast, HRTPP uncovers clinically meaningful rules, such as "if a high anion gap occurs before low heart rate, then death occurs", highlighting the link between metabolic disorders and circulatory failure. These rules enhance accuracy and align with medical logic, aiding clinical decision-making.
@@ 5.3 Rule Analysis
5.3.2 Rule Distribution Analysis.
Understanding mortality risk indicators in CAD patients is crucial for better management and prediction. To analyze these factors, we created a graph network based on medical guidelines [22], visually mapping their hierarchical relationships in Fig. 2. As shown in the Fig. 2, arterial blood pressure, heart rate, C-reactive protein and cholesterol directly effect mortality, while other indicators influence mortality through them. Hence, we classify the rules into direct or indirect rules according to their relationships. Table 3 shows direct indicators in green and indirect indicators in blue. HRTPP identifies the most direct rules, showing strong ability in capturing key CAD mortality factors. HRTPP also balances direct and indirect rules well, focusing on major factors without omitting minor details, thus enhancing its applicability in medical decision-making.
5.4 Stability Analysis
Model stability is crucial in medical decision support, ensuring consistency across conditions. To evaluate model stability, we recorded the frequency of four direct CAD indicators in five runs, using a five-fold cross-validation. Fig. 3 presents the results, where HR, ABP, Cho, and CRP denote Heart Rate, Arterial Blood Pressure, Cholesterol, and C-reactive Protein, respectively. As shown in the figure, HRTPP has the highest frequency of direct rules, which shows the best stability and consistency. Clinically, its stability ensures interpretable results by minimizing fluctuations caused by minor data changes, thereby enhancing decision-making accuracy and supporting personalized treatment.
HRABPChoCRPAvg.0%2020%4040%6060%8080%100100%8080%100100%4040%2020%6060%4040%6060%2020%0%3030%6060%100100%6060%8080%7575%100100%100100%8080%6060%8585%Medical indicationsFrequencyTELLERCLNNCLUSTERHRTPP
@@ 5.4 Stability Analysis
5.5 Case Study
In this section, we show the mechanism by which rules and events affect the target. When a rule is triggered, its strength increases instantaneously, reflecting the potential risk associated with the condition. Fig. 4 shows the intensity curve of a CAD patient. In this figure, abnormalities in physiological indicators triggered key rules over time. At 17.19h, high blood glucose appeared, followed by infection-related signs, increasing rule strength. At 60.26h, metabolic imbalance was triggered by an elevated anion gap and low heart rate. At 92.43h, high INR (International Normalized Ratio) and low heart rate indicated coagulation abnormalities. Finally, at 123.58h, decreased heart rate and increased blood pressure triggered the end-event rule, leading to a sharp rise in intensity and death.
Figure 4: Overall intensity over time for a CAD patient. The red curve represents the intensity of the target variable, reflecting the probability of death. Key events affecting the mortality are annotated below the curve, while the timing of logic rules is marked above. The entire figure illustrates the patient’s full clinical progression in the ICU.
This mechanism enables real-time monitoring and early warning of high-risk conditions while providing interpretable analysis of disease progression to support clinical decisions. By applying rule-guided dynamic analysis, we identify key causal relationships in CAD progression and detect turning points in disease deterioration, allowing clinicians to intervene precisely at critical moments. These rules are not just data-driven predictions but verifiable medical reasoning, ensuring transparency and reliability in CAD treatment.
5.6 Ablation Study
We conducted an ablation study to assess the impact of Numerical Feature Augmentation (NFA) on model performance. We classify models into two categories. Category I uses only basic intensity, like probabilistic TPP models. Category II uses basic and rule-based intensity. Table 4 shows the NLL of two categories. The introduction of NFA universally improve model performance across datasets, with the sole exception of AKI. The standard metrics employed in AKI evaluation underrepresent clinically significant intermediate events. This limitation is compounded by NFA focus on global rule coverage rather than domain-specific critical pattern recognition. These findings validate the effectiveness of numerical intensity in medical event modeling.
Table 4: Ablation study of NFA for NLL performance.
@@ 6 Conclusion
7 Acknowledgements
This work is supported by National Natural Science Foundation of China (62406270), Shanghai Rising-Star Program (24YF2748800).